# QWENDY: Gene Regulatory Network Inference by Quadruple Covariance Matrices

### v3

@@ Abstract
Knowing gene regulatory networks (GRNs) is important for understanding various biological mechanisms. In this paper, we present a method, QWENDY, that uses single-cell gene expression data measured at four time points to infer GRNs. Based on a linear gene expression model, it solves the transformation of the covariance matrices. Unlike its predecessor WENDY, QWENDY avoids solving a non-convex optimization problem and produces a unique solution. We test the performance of QWENDY on two experimental data sets and two synthetic data sets. Compared to previously tested methods on the same data sets, QWENDY ranks the first on experimental data, although it does not perform well on synthetic data.
@@ 1 Introduction
For such single-cell gene expression data measured at multiple time points, although it is the most informative data type under current technology, there are only a few GRN methods developed specifically for this data type [12]. We have developed the WENDY method to infer the GRN with this data type [13]. It needs measurement at two time points, and solves the transformation between the covariance matrices of two probability distributions, where the transformation is determined by the GRN. However, it needs to solve a non-convex optimization problem, which has infinitely many solutions, and WENDY will output one solution, determined by the numerical solver chosen in the realization.
@@
In a previous paper [14], we have tested 16 GRN inference methods on two experimental data sets and two synthetic data sets. In this paper, we test the performance of QWENDY on the same data sets and compare it with previous methods. In all 17 GRN inference methods, QWENDY ranks the first on experimental data, although it does not perform well on synthetic data. These results suggest that QWENDY is a promising GRN inference method that warrants further testing and refinement.
Section 2 reviews other GRN inference methods. Section 3 introduces the QWENDY method, including a proof of its correctness. Section 4 tests the performance of QWENDY on four data sets. We finish with some discussions in Section 5.
@@ 2 Literature review
Model-based methods construct concrete models for gene expression under regulation, and fit the expression data to the models to determine model parameters and reconstruct the GRN. Some methods [17] require measuring the same cell multiple times, which is not quite applicable for now. Some methods [18, 19] use average expression levels measured over many cells (bulk level), which do not utilize the rich information in the single-cell level measurements. Some methods [20] only work on single-cell data at one time point. For single-cell gene expression data measured at multiple time points, where each cell is measured only once, we only know one model-based method, WENDY. Therefore, we develop QWENDY as an alternative.
@@
Deep learning-based GRN inference methods [21, 22, 23, 24, 25] generally use neural networks as black boxes, without integrating them with gene expression models. Therefore, lacking of interpretability is a common problem.
@@
For more detailed summaries of deep learning-based GRN inference methods and approaches to enhance existing methods, readers may refer to the literature review section of the WENDY paper [13].
@@ 3.1 Setup
From this equation, WENDY directly solves $B$ (and thus $A$) by minimizing $||K_{1}-B^{\mathrm{T}}K_{0}B||_{\mathrm{F}}^{2}$, where $\mathrm{F}$ is the Frobenius norm. This problem is non-convex and has infinitely many solutions. WENDY outputs one solution of them.
@@
In this paper, we study whether we can better solve $B$ with data from more time points, especially $K_{2},K_{3}$. Since $B^{\mathrm{T}}KB=(-B)^{\mathrm{T}}K(-B)$, we cannot distinguish between $B$ and $-B$ from $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$.
@@ 3.2 Algorithm details
In this section, given general covariance matrices $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$ that may not satisfy Eqs. 13 for any $B_{0}$, we introduce a procedure to calculate $B$ that approximately solves Eqs. 13. Under some mild conditions (some matrices are invertible and have distinct eigenvalues), $B$ can be uniquely determined up to a $\pm$ sign. Then we use $\mathbf{x}_{0}$, $\mathbf{x}_{1}$, $\mathbf{x}_{2}$, $\mathbf{x}_{3}$ to distinguish between $B$ and $-B$. The whole procedure is named as the QWENDY method.
@@
$||L_{1}L_{1}^{\mathrm{T}}-L_{1}OO^{\mathrm{T}}L_{1}^{\mathrm{T}}||_{\mathrm{F}}^{2}.$
Therefore, Eq. 7 is minimized to 00 if and only if $O$ is orthonormal: $OO^{\mathrm{T}}=I$.
@@
$||L_{1}^{-1}(K_{2}-B^{\mathrm{T}}K_{1}B)L_{1}^{-\mathrm{T}}||_{\mathrm{F}}^{2}$ (8)
@@
$||P_{2}D_{2}P_{2}^{\mathrm{T}}-OP_{1}D_{1}P_{1}^{\mathrm{T}}O^{\mathrm{T}}||^{2}_{\mathrm{F}}=||D_{2}-P_{2}^{\mathrm{T}}OP_{1}D_{1}P_{1}^{\mathrm{T}}O^{\mathrm{T}}P_{2}||^{2}_{\mathrm{F}},$
@@
For diagonal $D_{1},D_{2}$ with strictly increasing diagonal elements and any orthonormal $W$, Eq. 9 is minimized when $W$ is diagonal, and the diagonal elements are ±$\pm 1$. (There are $2^{n}$ possibilities for such $W$.)
@@
$\begin{split}&||D_{2}-WD_{1}W^{\mathrm{T}}||_{\mathrm{F}}^{2}=||D_{2}||_{\mathrm{F}}^{2}+||WD_{1}W^{\mathrm{T}}||_{\mathrm{F}}^{2}-2\sum_{i=1}^{n}[D_{2}\otimes(WD_{1}W^{\mathrm{T}})]_{ii}\\ =&||D_{2}||_{\mathrm{F}}^{2}+||D_{1}||_{\mathrm{F}}^{2}-2\sum_{i=1}^{n}d_{2,i}\sum_{j=1}^{n}W_{ij}^{2}d_{1,j}\\ =&||D_{2}||_{\mathrm{F}}^{2}+||D_{1}||_{\mathrm{F}}^{2}-2\mathbf{d}_{1}(W\otimes W)\mathbf{d}_{2}^{\mathrm{T}},\end{split}$
@@
meaning that $W$ is diagonal, and diagonal elements are ±$\pm 1$. ∎
@@
$||L_{1}^{-1}(K_{3}-B^{\mathrm{T}}K_{2}B)L_{1}^{-\mathrm{T}}||_{\mathrm{F}}^{2}.$ (10)
@@
$\begin{split}&||P_{2}^{\mathrm{T}}L_{1}^{-1}K_{3}L_{1}^{-\mathrm{T}}P_{2}-WP_{1}^{\mathrm{T}}L_{0}^{-1}K_{2}L_{0}^{-\mathrm{T}}P_{1}W||_{\mathrm{F}}^{2}\\ =&||G-WHW||_{\mathrm{F}}^{2}=||G||_{\mathrm{F}}^{2}+||WHW||_{\mathrm{F}}^{2}-2||G\otimes(WHW)||_{\mathrm{F}}^{2}\\ =&||G||_{\mathrm{F}}^{2}+||H||_{\mathrm{F}}^{2}-2||G\otimes(WHW)||_{\mathrm{F}}^{2}.\end{split}$ (11)
We want to find diagonal $W$ with ±$\pm 1$ that minimizes Eq. 10, which is equivalent to maximizing $||G\otimes(WHW)||_{\mathrm{F}}^{2}$. Define $C=G\otimes H$, which is still positive definite and symmetric by Schur product theorem [42]. Assume that $C$ does not have repeated eigenvalues. Denote the diagonal elements of $W$ by $\mathbf{w}=[w_{1},\ldots,w_{n}]$. Then we need to maximize
$||G\otimes(WHW)||_{\mathrm{F}}^{2}=\mathbf{w}C\mathbf{w}^{\mathrm{T}}.$
Now we relax this problem from $\mathbf{w}$ with $w_{i}=\pm 1$ to general $\mathbf{v}$ with $||\mathbf{v}||_{2}^{2}=n$:
$\max_{||\mathbf{v}||_{2}^{2}=n}\mathbf{v}C\mathbf{v}^{\mathrm{T}}.$
@@
After obtaining $\mathbf{v}$, we just need to project it to $[w_{1},\ldots,w_{n}]$ with $w_{i}=\pm 1$ by taking the sign of each term: $w_{i}=\mathrm{sign}(v_{i})$. Then construct $W$ by putting $[w_{1},\ldots,w_{n}]$ on the diagonal.
@@
Step (4): For $B$ and $-B$ from Step (3), we determine which satisfies Eqs. 46 better. Notice that Eqs. 46 share the same unknown $\mathbf{c}$. Define
@@
$||\mathbf{c}_{0}-\bar{\mathbf{c}}||_{2}^{2}+||\mathbf{c}_{1}-\bar{\mathbf{c}}||_{2}^{2}+||\mathbf{c}_{2}-\bar{\mathbf{c}}||_{2}^{2}=||\mathbf{c}_{0}||_{2}^{2}+||\mathbf{c}_{1}||_{2}^{2}+||\mathbf{c}_{2}||_{2}^{2}-3||\bar{\mathbf{c}}||_{2}^{2}.$
We just need to compare the errors of $B$ and $-B$ and choose the smaller one as the output.
With expression data from four time points, we uniquely determine $B$. See Algorithm 1 for the workflow of the QWENDY method.
@@
    Input expression levels of $n$ genes over $m$ cells at time points 00, $t$, $2t$, $3t$
@@
    Compare total squared errors of $B$ and $-B$ for Eqs. 46
@@
    Output $B$ or $-B$, the one that corresponds to the smaller error, and the GRN $A=(B-I)/t$
@@ 3.3 Correctness of the QWENDY method
If covariance matrices $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$ and mean levels $\mathbf{x}_{0}$, $\mathbf{x}_{1}$, $\mathbf{x}_{2}$, $\mathbf{x}_{3}$ satisfy Eqs. 16 for some $B_{0}$, then QWENDY will output $B_{0}$.
@@
If $K_{1}=B_{0}^{\mathrm{T}}K_{0}B_{0}$, then
@@
If $K_{2}=B_{0}^{\mathrm{T}}K_{1}B_{0}$, then
@@
Since $P_{2}^{\mathrm{T}}O_{0}P_{1}$ is orthonormal, $D_{1}$ and $D_{2}$ are similar and have the same eigenvalues. Therefore, $D_{1}$ and $D_{2}$ as diagonal matrices with increasing diagonal values are equal. Assume that the diagonal elements $d_{1},\ldots,d_{n}$ of $D_{1}$ are distinct. Define $W_{0}=P_{2}^{\mathrm{T}}O_{0}P_{1}$, then
@@
and $W_{0}[i,j]d_{i}=W_{0}[i,j]d_{j}$. For $i\neq j$, we have $d_{i}\neq d_{j}$, meaning that $W_{0}[i,j]=0$. Now $W_{0}$ is diagonal and orthonormal, implying that its diagonal elements are $1$ or $-1$, and
@@
Define $\mathbf{w}_{0}=[w_{1},\ldots,w_{n}]$ to be the diagonal elements of $W_{0}$. Then from Step (3), $W_{0}$ minimizes Eq. 10, and $\mathbf{w}_{0}$ is the solution to
$\max_{||\mathbf{v}||_{2}^{2}=n}\mathbf{v}C\mathbf{v}^{\mathrm{T}},$
namely the eigenvector of the largest eigenvalue. Since $w_{i}=\pm 1$, the projection $w_{i}=\mathrm{sign}(v_{i})$ has no effect, and
@@
Therefore, $-B_{0}$ does not satisfy Eqs. 46, and Step (4) chooses the correct $B_{0}$ from the two possibilities in Step (3). ∎
@@ 4 Performance on synthetic and experimental data
The same as in the WENDY paper and the TRENDY paper [13, 14], we test the QWENDY method on two synthetic data sets and two experimental data sets. SINC data set [14] and DREAM4 data set [43] are generated by simulating stochastic differential equation systems. THP-1 data set is from monocytic THP-1 human myeloid leukemia cells [44]. hESC data set is from human embryonic stem cell-derived progenitor cells[45]. THP-1 and hESC data sets each has only one group of data (one ground truth GRN and the corresponding expression levels); DREAM4 data set has five groups of data; SINC data set has 1000 groups of data.
For SINC data set and DREAM4 data set, we use data from any four consecutive time points and take average. For THP-1 data set and hESC data set, we choose any four time points with equal difference. This is due to the limitation that QWENDY is derived for four evenly spaced time points.
We test the performance of the QWENDY method on these four data sets and compare it with the performance of 16 methods tested on the same data sets in the TRENDY paper [14]. We compare the inferred GRN and the ground truth GRN by calculating the AUROC and AUPRC scores [13]. These two scores, both between 00 and $1$, evaluate the level of matching, where $1$ means perfect match, and 00 means perfect mismatch.

### v4

@@ Abstract
Knowing gene regulatory networks (GRNs) is important for understanding various biological mechanisms. In this paper, we present a method, QWENDY, that uses single-cell gene expression data measured at four time points to infer GRNs. Based on a linear gene expression model, it solves the transformation of the covariance matrices. Unlike its predecessor WENDY, QWENDY avoids solving a non-convex optimization problem and produces a unique solution. We test the performance of QWENDY on three experimental data sets and two synthetic data sets. Compared to previously tested methods on the same data sets, QWENDY ranks the first on experimental data, although it does not perform well on synthetic data.
@@ 1 Introduction
For such single-cell gene expression data measured at multiple time points, although it is the most informative data type under current technology, there are only a few GRN methods developed specifically for this data type [12, 13]. We have developed the WENDY method to infer the GRN with this data type [14]. It needs measurement at two time points, and solves the transformation between the covariance matrices of two probability distributions, where the transformation is determined by the GRN. However, it needs to solve a non-convex optimization problem, which has infinitely many solutions, and WENDY will output one solution, determined by the numerical solver chosen in the realization.
@@
In a previous paper [15], we have tested 16 GRN inference methods on two experimental data sets and two synthetic data sets. In this paper, we test the performance of QWENDY on the same data sets and an extra experimental data set, and compare QWENDY with previous methods. In all 17 GRN inference methods, QWENDY ranks the first on experimental data, although it does not perform well on synthetic data. These results suggest that QWENDY is a promising GRN inference method that warrants further testing and refinement.
Section 2 reviews other GRN inference methods. Section 3 introduces the QWENDY method, including a proof of its correctness. Section 4 tests the performance of QWENDY on four data sets. After conclusions in Section 5, we finish with some discussions in Section 6.
@@ 2 Literature review
Model-based methods construct concrete models for gene expression under regulation, and fit the expression data to the models to determine model parameters and reconstruct the GRN. Some methods [18] require measuring the same cell multiple times, which is not quite applicable for now. Some methods [19, 20] use average expression levels measured over many cells (bulk level), which do not utilize the rich information in the single-cell level measurements. Some methods [21, 22] only work on single-cell data at one time point. For single-cell gene expression data measured at multiple time points, where each cell is measured only once, we only know one model-based method, WENDY. Therefore, we develop QWENDY as an alternative.
@@
Deep learning-based GRN inference methods [23, 24, 25, 13, 26, 27] generally use neural networks as black boxes, without integrating them with gene expression models. Therefore, lacking of interpretability is a common problem.
@@
For more detailed summaries of deep learning-based GRN inference methods and approaches to enhance existing methods, readers may refer to the literature review section of the TRENDY paper [15].
@@
Gene expression and regulation can have spatial patterns. There are some studies on GRN inference from spatial transcriptomics data [42, 43, 44, 45, 46, 47]. We shall assume spatial homogeneity in this paper.
@@ 3.1 Setup
From this equation, WENDY directly solves $B$ (and thus $A$) by minimizing $K_{1}-B^{\mathrm{T}}K_{0}B||_{\mathrm{F}}^{2}$, where $\mathrm{F}$ is the Frobenius norm. This problem is non-convex and has infinitely many solutions. WENDY outputs one solution of them.
@@
In this paper, we study whether we can better solve $B$ with data from more time points, especially $K_{2},K_{3}$. Since $B^{\mathrm{T}}KB=(-B)^{\mathrm{T}}K(-B)$, we cannot distinguish between $B$ and B-B from $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$.
@@ 3.2 Algorithm details
In this section, given general covariance matrices $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$ that may not satisfy Eqs. 13 for any $B_{0}$, we introduce a procedure to calculate $B$ that approximately solves Eqs. 13. Under some mild conditions (some matrices are invertible and have distinct eigenvalues), $B$ can be uniquely determined up to a $\pm$ sign. Then we use $\mathbf{x}_{0}$, $\mathbf{x}_{1}$, $\mathbf{x}_{2}$, $\mathbf{x}_{3}$ to distinguish between $B$ and B-B. The whole procedure is named as the QWENDY method.
@@
$L_{1}L_{1}^{\mathrm{T}}-L_{1}OO^{\mathrm{T}}L_{1}^{\mathrm{T}}||_{\mathrm{F}}^{2}.$
Therefore, Eq. 7 is minimized to 0 if and only if $O$ is orthonormal: $OO^{\mathrm{T}}=I$.
@@
$L_{1}^{-1}(K_{2}-B^{\mathrm{T}}K_{1}B)L_{1}^{-\mathrm{T}}||_{\mathrm{F}}^{2}$ (8)
@@
[…]
@@
For diagonal $D_{1},D_{2}$ with strictly increasing diagonal elements and any orthonormal $W$, Eq. 9 is minimized when $W$ is diagonal, and the diagonal elements are $\pm 1$. (There are $2^{n}$ possibilities for such $W$.)
@@
[…]
@@
meaning that $W$ is diagonal, and diagonal elements are $\pm 1$. ∎
@@
$L_{1}^{-1}(K_{3}-B^{\mathrm{T}}K_{2}B)L_{1}^{-\mathrm{T}}||_{\mathrm{F}}^{2}.$ (10)
@@
[…]
We want to find diagonal $W$ with $\pm 1$ that minimizes Eq. 10, which is equivalent to maximizing $G\otimes(WHW)||_{\mathrm{F}}^{2}$. Define $C=G\otimes H$, which is still positive definite and symmetric by Schur product theorem [50]. Assume that $C$ does not have repeated eigenvalues. Denote the diagonal elements of $W$ by $\mathbf{w}=[w_{1},\ldots,w_{n}]$. Then we need to maximize
$G\otimes(WHW)||_{\mathrm{F}}^{2}=\mathbf{w}C\mathbf{w}^{\mathrm{T}}.$
Now we relax this problem from $\mathbf{w}$ with $w_{i}=\pm 1$ to general $\mathbf{v}$ with $\mathbf{v}||_{2}^{2}=n$:
[…]
@@
After obtaining $\mathbf{v}$, we just need to project it to $w_{1},\ldots,w_{n}]$ with $w_{i}=\pm 1$ by taking the sign of each term: $w_{i}=\mathrm{sign}(v_{i})$. Then construct $W$ by putting $w_{1},\ldots,w_{n}]$ on the diagonal.
@@
Step (4): For $B$ and B-B from Step (3), we determine which satisfies Eqs. 46 better. Notice that Eqs. 46 share the same unknown $\mathbf{c}$. Define
@@
[…]
We just need to compare the errors of $B$ and B-B and choose the smaller one as the output.
With expression data from four time points, we uniquely determine $B$. See Algorithm 1 for the workflow of the QWENDY method. The assumptions in the derivation should be checked.
@@
    Input expression levels of $n$ genes over $m$ cells at time points 0, $t$, $2t$, $3t$
@@
    (Check $L_{0}^{-1}K_{1}L_{0}^{-\mathrm{T}}$ and $L_{1}^{-1}K_{2}L_{1}^{-\mathrm{T}}$ each has distinct eigenvalues)
@@
    Compare total squared errors of $B$ and B-B for Eqs. 46
@@
    Output $B$ or B-B, the one that corresponds to the smaller error, and the GRN A=(BI)/tA=(B-I)/t
    (If any check fails, output a warning that the result might be problematic)
@@ 3.3 Correctness of the QWENDY method
If covariance matrices $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$ and mean levels $\mathbf{x}_{0}$, $\mathbf{x}_{1}$, $\mathbf{x}_{2}$, $\mathbf{x}_{3}$ satisfy Eqs. 16 for some $B_{0}$, then QWENDY will output $B_{0}$ under the following conditions:
(1) $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$ are invertible;
(2) $L_{0}^{-1}K_{1}L_{0}^{-\mathrm{T}}$ and $L_{1}^{-1}K_{2}L_{1}^{-\mathrm{T}}$ each has distinct eigenvalues;
(3) $C$ has distinct eigenvalues.
@@
Due to condition (1), $L_{0}^{-1}$ and $L_{1}^{-1}$ exist. If $K_{1}=B_{0}^{\mathrm{T}}K_{0}B_{0}$, then
@@
Due to condition (2), $P_{1},D_{1},P_{2},D_{2}$ are uniquely defined. If $K_{2}=B_{0}^{\mathrm{T}}K_{1}B_{0}$, then
@@
Since $P_{2}^{\mathrm{T}}O_{0}P_{1}$ is orthonormal, $D_{1}$ and $D_{2}$ are similar and have the same eigenvalues. Therefore, $D_{1}$ and $D_{2}$ as diagonal matrices with increasing diagonal values are equal. Due to condition (2), $d_{1},\ldots,d_{n}$ of $D_{1}$ are distinct. Define $W_{0}=P_{2}^{\mathrm{T}}O_{0}P_{1}$, then
@@
and $W_{0}[i,j]d_{i}=W_{0}[i,j]d_{j}$. For $i\neq j$, we have $d_{i}\neq d_{j}$, meaning that $W_{0}[i,j]=0$. Now $W_{0}$ is diagonal and orthonormal, implying that its diagonal elements are $1$ or 1-1, and
@@
Define $\mathbf{w}_{0}=[w_{1},\ldots,w_{n}]$ to be the diagonal elements of $W_{0}$. Then from Step (3), $W_{0}$ minimizes Eq. 10, and $\mathbf{w}_{0}$ is the unique solution to
[…]
namely the eigenvector of the largest eigenvalue. Here the uniqueness of the solution is from condition (3). Since $w_{i}=\pm 1$, the projection $w_{i}=\mathrm{sign}(v_{i})$ has no effect, and
@@
Therefore, $B_{0}$ does not satisfy Eqs. 46, and Step (4) chooses the correct $B_{0}$ from the two possibilities in Step (3). ∎
@@ 4 Performance on synthetic and experimental data
The same as in the WENDY paper and the TRENDY paper [14, 15], we test the QWENDY method on synthetic data sets SINC and DREAM4, and experimental data sets THP-1 and hESC. SINC data set [15] and DREAM4 data set [51] are generated by simulating stochastic differential equation systems. THP-1 data set is from monocytic THP-1 human myeloid leukemia cells [52]. hESC data set is from human embryonic stem cell-derived progenitor cells[53]. Besides, we also test the QWENDY method on experimental data set mESC [54], which is from primitive endoderm cells differentiated from mouse embryonic stem cells. THP-1, hESC, and mESC data sets each has only one group of data (one ground truth GRN and the corresponding expression levels); DREAM4 data set has five groups of data; SINC data set has 1000 groups of data.
For SINC data set and DREAM4 data set, we use data from any four consecutive time points and take average. For THP-1, hESC, and mESC data sets, we choose any four time points with equal difference. This is due to the limitation that QWENDY is derived for four evenly spaced time points.
We test the performance of the QWENDY method on these five data sets and compare it with the performance of 16 methods tested on the same data sets in the TRENDY paper [15]. We compare the inferred GRN and the ground truth GRN by calculating the AUROC and AUPRC scores [14]. These two scores, both between 0 and $1$, evaluate the level of matching under different thresholds, where $1$ means perfect match, and 0 means perfect mismatch. AUROC and AUPRC have an advantage that they do not have parameters that can be chosen manually, which guarantees fair comparison.